Status Report

SOHO Daily Meeting Minutes for Wednesday, 27 February 2002

By SpaceRef Editor
February 27, 2002
Filed under , ,

* HESSI target for today: 9830
HESSI target for tomorrow: the active region complex
containing NOAA 9839, 9841, 9842, and 9830

* Campaign number 6850 has been created for flagging studies
in full or partial support of the HESSI default collaboration.

* CMS testing this morning was successful

* The next SSR dump will be Thursday

* The visit from the members of the House Appropriations
Committee staff has been postponed.

* Solar science seminar Wed Feb 27, 2002 at 3:30pm in the EAF by
David Rust (JHU/APL) on Magnetic Flux Emergence in AR8844 – Flare
Genesis Observations.

* CELIAS patch uplink will be at 17 UT today

Spacecraft Status: Nominal

Spacecraft Anomalies:
057/2343 CEPAC ESU Data Request Error; recovered
autonomously after 4 minutes.

Accomplished Activities: VIRGO, RSL

Planned Activities: VIRGO, SWAN

Upcoming Operations: None

Ground Anomalies:
057/1831 D27 AOS was 22 minutes late due to problems with
the antenna controller console. No data loss as
D66 was able to extend their support. MDI
continuous campaign interrupted with an unscheduled
transition to IDLE.
058/0953 Lost 1 minute 33 seconds of unrecoverable TLM as
D66’s antennna went to brake.

02/26 DOY 057 – 02/27 DOY 058

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. AR observations and polarity reversal in the north.
Tomorrow: Similar, with polarity reversal in the south.

UVCS: Nominal. Long synoptic. Tomorrow: Streamer study and synoptic.

LASCO: Nominal. C2/C3 synoptics.

EIT: Nominal. 195 Angstrom CME watch and synoptics.

MDI: Nominal. Full disk magnetograms and dopplergrams for continuous

TRACE: Nominal. Observing 9845.

SpaceRef staff editor.